Four years ago,Clara was 6 times as old as Mara. Four years from now, Clara will only be twice as old as Mara. How old are Mara and Clara this time?
Answer by stanbon(75887)

Four years ago,Clara was 6 times as old as Mara. Four years from now, Clara will only be twice as old as Mara. How old are Mara and Clara this time?
-------------------
Equations:
(c-4) = 6(m-6)
(c+4) = 2(m+4)
------
Rearrange:
c - 6m = -32
c - 2m = 4
----
Subtract and solve for m:
4m = 36
m = 9 yrs (Mara's age now)
----
Solve for "c"
c = 2m+4
c = 2*9+4
c 22 yrs (Clara's age now)
=========================
Cheers,
Stan H.
============